If the oxygen gas sample in#1 which is "a sample of oxygen gas,which is saturated with water vapor,is kept in a 10 L vessel at 30°C and has a pressure of 758 Torr. If the pressure of the water vapour at this temperature is 31.8 what would be the pressure of the dry oxygen" passed through a drier that decreased the pressure of the mixture to 750 torr and the pressure exerted by the water vapour is only 80 percent of the saturated vapour pressure at the given temperature, what would be the pressure of the dry oxygen?


1
Expert's answer
2021-04-12T01:18:16-0400

p(total)=p1+p2.

p2=p(total)-p1 = 758-31.8 = 726.2 torr.